# ComGPT: Detecting Local Community Structure with Large Language Models

### v2

@@ Introduction
Five real-world datasets, including Dolphins (Lusseau 2003), Football (Girvan and Newman 2002), Polbooks (Krebs 2004), Amazon (Yang and Leskovec 2012), and DBLP (Yang and Leskovec 2012) are utilized to test the performance of algorithms. The Dolphins network is based on observations of associations between pairs of dolphins, where edges represent associations between dolphins. It contains 62 nodes and 159 edges. The Football network contains 115 nodes and 613 edges. Nodes represent college football teams, communities correspond to different leagues, and edges correspond to games played between them. The Polbooks network is a co-purchasing network for books. Each node in the dataset represents a book about US politics. An edge between two books indicates that they are often purchased together by customers. It contains 105 nodes and 441 edges. The Amazon network represents product co-purchasing relationships, consisting of 334,863 nodes and 925,872 edges. Nodes representing products and edges indicate which products are frequently co-purchased. The DBLP network is a co-authorship network comprising 317,080 nodes and 1,049,866 edges. Here, nodes represent authors and edges indicate that two authors have co-authored a paper.
Baselines
@@
To evaluate the performance of ComGPT, we compare it with M method (Luo, Wang, and Promislow 2008), R method (Clauset 2005), DMF_M (Luo et al. 2018), DMF_R (Luo et al. 2018), LCDNN (Luo et al. 2020), and LS (Shi et al. 2024).
Implementation Details
@@
ComGPT uses gpt-3.5-turbo-0125. The number of potential nodes $k$ is set to 10. All experiments are conducted under Windows 10 operating system (CPU: 11th Gen Intel(R) Core(TM) i7-11700K @ 3.60GHZ, Memory: 32 GB.). Due to the monetary costs of GPT-3.5-turbo-0125, we execute ComGPT twice for each starting seed node to handle GPT hallucinations.
Evaluation metrics
@@
The evaluation metrics Fscore (Powers 2011) and Jaccard (Jaccard 1912) are used to evaluate the detected communities. For an overlapped seed node, all the communities it belongs to are integrated as a real local community. Fscore is the harmonic mean of precision and recall. The values of Fscore and Jaccard are between 0 and 1. The larger value implies a better algorithm performance.
Experimental results
Results on networks
@@
For the Dolphins, Polbooks, and Football datasets, all node are selected as seed nodes. For the Amazon and DBLP datasets, we randomly select 200 nodes from the top 5,000 communities as seed nodes. The average Fscore and Jaccard of all seed nodes is calculated. The detailed experimental results are shown in Table 1.
@@
Table 1 shows that ComGPT outperforms the M method, R method, DMF_M, DMF_R, LCDNN, and LS on most datasets. Among all comparison methods, ComGPT exhibits the best performance on the Amazon and DBLP datasets and ranks second or third on the other three datasets. Although LS achieves the best performance on the Dolphins and Polbooks datasets, its performance on the remaining three datasets is the lowest among all comparison methods. Overall, ComGPT is superior to LS. ComGPT is superior to other methods because GPT-3.5-turbo has excellent comprehension to find the suitable node based on the NSG prompt. On the Football dataset, ComGPT performs worse than DMF_M, DMF_R, and LCDNN. It is because Football contains several tightly interconnected communities, which leads GPT-3.5-turbo to erroneously identify them as a single community. Except for the DBLP dataset, DMF_M, DMF_R, and LCDNN perform better than the M method and R method. The reason why DMF_M, DMF_R, and LCDNN outperform the M method and R method on most datasets is that the latter two methods maximize local modularity ($M$ and $R$), making them sensitive to the choice of seed nodes, known as seed-dependent problem. In contrast, DMF_M, DMF_R, and LCDNN implement various strategies to alleviate this problem. Specifically, DMF_M, and DMF_R, utilize membership functions to select nodes with large degrees. LCDNN expands the community based on the fuzzy relations between nodes and NGC nodes.
Ablation study
Method GPT SK NSG NSU Fscore Jaccard
1 NO NO NO NO 0.4666 0.3352
2 YES NO YES YES 0.3711 0.2414
3 YES YES NO YES 0.4855 0.3299
4 YES YES YES NO 0.7134 0.6307
ComGPT YES YES YES YES 0.7688 0.7009
Table 2: Results of ablation experiment on the Dolphins dataset.
@@
To explore the necessity of GPT-3.5-turbo, SK (Supplementary knowledge in graph coding), NSG, and NSU prompts in ComGPT, we conduct ablation experiments on the Dolphins dataset with 50 randomly selected seed nodes. Four simplified versions of ComGPT are constructed as follows: 1) Method 1 without GPT-3.5-turbo. The first node in $PN$ is selected for community expansion until $PN$ is empty; 2) Method 2 without SK; 3) Method 3 without NSG prompt; and 4) Method 4 without NSU prompt. When calculating Fscore and Jaccard, it is necessary to obtain the intersection between the detected community, which may include duplicate nodes, and the true community. Here, duplicate nodes in the detected community are regarded as a single node. To prevent cyclic selection of duplicate nodes, community expansion is terminated once the number of duplicate nodes exceeds one-third of the number of nodes in the community.
@@
Table 2 shows the experimental results of Methods 1, 2, 3, 4, and ComGPT. Overall, ComGPT outperforms Methods 1, 2, 3, and 4, indicating that GPT-3.5-turbo, SK, NSG, and NSU prompt play a crucial role in ComGPT, respectively. ComGPT is better than Method 1 in terms of Fscore and Jaccard, which demonstrates the effectiveness and importance of GPT-3.5-turbo. The reason ComGPT outperforms Method 2 is that its graph coding incorporates community knowledge. Compared with graph encoding methods that contain only the graph topology, supplementing community knowledge enables the GPT-3.5-turbo to better understand the community structure and thus make more accurate judgments. ComGPT also outperforms Method 3 and 4, suggesting that NSG and NSU prompts can enhance its performance in local community detection.
Prompt comparisons
@@
Table 3: Description of the prompts.
 Prompts Dolphins Polbooks Football
Zero-shot Jaccard 0.3299 0.2191 0.5829
Fscore 0.4855 0.3490 0.6774
Few-shot Jaccard 0.3584 0.2703 0.7145
Fscore 0.5128 0.4078 0.7764
CoT Jaccard 0.3250 0.2078 0.6168
Fscore 0.4785 0.3347 0.7196
BaG Jaccard 0.2772 0.1813 0.6050
Fscore 0.4235 0.2988 0.7131
NSG Jaccard 0.7009 0.6624 0.8045
Fscore 0.7688 0.7418 0.8555
Table 4: Results of different prompts.
To test the performance of the NSG prompt within ComGPT, we replaced the NSG prompt with four typical prompt approaches: Zero-shot (Brown et al. 2020), Few-shot (Brown et al. 2020), CoT (Wei et al. 2022), and BaG (Wang et al. 2023) prompts, as shown in Table 3. These alternate versions of the algorithm serve as baseline models for comparison. We conducted comparative experiments by randomly selecting 50 seed nodes in the Polbooks, Dolphins, and Football datasets. The detected communities may contain duplicate nodes, and the metrics values are calculated in the same way as in the ablation study.
Table 4 shows that the performance of ComGPT with NSG is better than that of ComGPT with Zero-shot, Few-shot, CoT, and BaG prompts. Specifically, on Football dataset, the ComGPT with CoT and BaG outperform the ComGPT with Zero-shot. The ComGPT with CoT and BaG is worse than ComGPT with Zero-shot on Dolphins and Polbooks datasets. Advanced prompts may have limited performance improvement or even counterproductive effects on complex graph reasoning tasks. For instance, advanced prompts like CoT increase the number of reasoning steps for GPT-3.5-turbo, which potentially raises the likelihood of errors in the final reasoning conclusion due to mistakes in one of the intermediate steps. The reason ComGPT with Few-shot outperforms ComGPT with Zero-shot on most datasets is that we provided some examples of how the task should be executed. These examples guide the model by showing the correct process needed to solve the task. On Dolphins, Polbooks, and Football datasets, ComGPT with NSG outperforms ComGPT with other prompts. This is because the design of CoT and BaG is oriented towards universal graph reasoning problems so that no relevant guidance or knowledge is utilized in dealing with local community detection. NSG is oriented to the field of local community detection and utilizes the guidance in local community detection.
Effect of the number of potential nodes on ComGPT
@@
We explore the effect of the number of potential nodes, denoted by $k$, on ComGPT on the Football dataset, with 50 seed nodes selected randomly. The values of $k$ are 5, 7, 10, 13, and 15. Figure 4 shows the performance of ComGPT under different $k$. As the $k$ increases, the performance of ComGPT remains generally stable, though with some fluctuations. Fluctuations are relatively low across different values, except for $k$ = 10, which exhibits higher variability due to the randomness of GPT-3.5-turbo output. To confirm this, we conduct repeated experiments for $k$ = 10 and the results are nearly identical to those obtained with other values of $k$. The above analysis suggests that the performance of ComGPT is less affected by the parameter $k$.
Conclusion

### v3

@@ III-A1 Datasets
The Dolphins network is based on observations of associations between pairs of dolphins, where edges represent associations between dolphins. The Football network is based on college football matches. Nodes represent college football teams, communities correspond to different leagues, and edges correspond to games played between them. The Polbooks network is a co-purchasing network for books. Each node in the dataset represents a book about US politics. An edge between two books indicates that they are often purchased together by customers. The Amazon network represents product co-purchasing relationships. Nodes representing products and edges indicate which products are frequently co-purchased. The DBLP network is a co-authorship network. Here, nodes represent authors and edges indicate that two authors have co-authored a paper.
III-A2 Baselines
@@
To evaluate the performance of ComGPT, we compare it with M method [4], R method [24], Louvain [31], DMF_M [29], DMF_R [29], Leiden [32], LCDNN [15], LS [33], and FLMIG [34], introduced as follows:
  1. 1.
    M method [4] and R method [24]. The M method and R method are greedy algorithms based on modularity. The given node constitutes the initial local community at first. M method (R method) iteratively selects and adds the node that maximizes the modularity MM($R$) to the local community until no node will decrease modularity MM($R$).
  2. 2.
    Louvain [31]. The Louvain first treats each node as an individual community and iteratively adjusts node assignments to maximize modularity until stability is reached. Then, it aggregates communities into new nodes, constructs a simplified graph, and repeats the optimization process until modularity converges.
  3. 3.
    DMF_M and DMF_R[29]. DMF_M and DMF_R analyze the formation process of the local community and divide it into three stages. Each stage corresponds to a different dynamic membership function.
  4. 4.
    Leiden [32]. Leiden is a derivative version of Louvain. Unlike Louvain, Leiden ensures well-connected communities through a refinement step. It then contracts communities into new nodes, constructs a simplified graph, and repeats the process until convergence, achieving more stable partitions.
  5. 5.
    LCDNN [15]. LCDNN relies on NGC nodes. The nodes are added to the local community based on the fuzzy relations between them and their NGC nodes. At each time, if the largest fuzzy relation is no less than half of the average fuzzy relation of the current local network, one node with the largest fuzzy relation is added to the local community.
  6. 6.
    LS [33]. LS is built on the notion of local dominance, where low-degree nodes are assigned to the basin of influence of high-degree nodes. Utilizing this principle, LS identifies community centers based on node degree and other local information. Finally, it divides communities according to the identified community centers.
  7. 7.
    FLMIG [34]. FLMIG is also a derivative version of Louvain. FLMIG enhances the Louvain Prune heuristic using an iterated greedy (IG) framework to maximize modularity in non-overlapping communities.
For baselines that require parameter settings, we select the optimal parameter within the given range as the final value. For Louvain, the parameter that controls the size of the communities takes values within [0.3, 0.5, 0.8, 1, 1.5]. For Leiden, the maximum total size of nodes in a community is tuned within [0, 50, 100, 150, 200]. For FLMIG, the parameters are set as explicitly provided by the authors.
III-A3 Implementation Details
@@
ComGPT uses gpt-3.5-turbo-0125. The number of potential nodes $k$ is set to 10. All experiments are conducted under Windows 10 operating system (CPU: 11th Gen Intel(R) Core(TM) i7-11700K @ 3.60GHZ, Memory: 32 GB.). Due to the monetary costs of GPT-3.5-turbo-0125, we execute ComGPT twice for each starting seed node to handle GPT hallucinations. The code is released on the author’s website [^1] https://github.com/cassius0301/ComGPT.
III-A4 Evaluation metrics
The evaluation metrics Jaccard [35] and Fscore [36] are used to evaluate the detected communities. Their formulas are as follows.
$\text{Jaccard}=\frac{\left|C_{\text{Found}}\cap C_{\text{True}}\right|}{\left|C_{\text{Found}}\cup C_{\text{True}}\right|},$ (2)
$\text{Fscore}=\frac{2\cdot\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}},$ (3)
$\text{Recall}=\frac{\left|C_{\text{Found}}\cap C_{\text{True}}\right|}{\left|C_{\text{True}}\right|},$ (4)
$\text{Precision}=\frac{\left|C_{\text{Found}}\cap C_{\text{True}}\right|}{\left|C_{\text{Found}}\right|},$ (5)
@@ III-A4 Evaluation metrics
where $C_{\text{Found}}$ is the local community that the algorithm discovers, and $C_{\text{True}}$ is the real local community to which the given starting node belongs. The values of Fscore and Jaccard are between 0 and 1. The larger value implies a better algorithm performance. For an overlapped seed node, all the communities it belongs to are integrated as a real local community.
III-B Experimental results
TABLE III: Experimental results on five datasets. On each dataset, the highest Fscore (Jaccard) is highlighted in bold, the second highest one is highlighted by underline, and the third highest one is highlighted with a superscript asterisk.
Dataset Metrics M R Louvain DMF_M DMF_R Leiden LCDNN LS FLMIG ComGPT
Dolphins Jaccard 0.3370 0.3128 0.4705 0.6270 0.6320 0.5307 0.6438* 0.7735 0.5359 0.7035
Fscore 0.4673 0.4478 0.5872 0.7254 0.7241 0.6320 0.7558* 0.8385 0.6517 0.7684
Polbooks Jaccard 0.4266 0.3986 0.6403 0.6449 0.6437 0.6374 0.6673 0.6919 0.6336 0.6468*
Fscore 0.5184 0.4960 0.7229 0.7294 0.7285 0.7237 0.7463 0.7764 0.7191 0.7320*
Football Jaccard 0.6117 0.6134 0.8516* 0.8546 0.8579 0.7762 0.7547 0.2406 0.5596 0.7571
Fscore 0.6706 0.6699 0.8890 0.8881* 0.8893 0.8361 0.8255 0.3593 0.6798 0.8103
Amazon Jaccard 0.6961 0.6771 0.1917 0.7633 0.7686 0.4741 0.7665* 0.6506 0.0479 0.7734
Fscore 0.7706 0.7554 0.2499 0.8353 0.8361* 0.5940 0.8372 0.7420 0.0642 0.8421
DBLP Jaccard 0.3239 0.3207* 0.1786 0.2945 0.2928 0.2076 0.2208 0.0650 0.0074 0.3259
Fscore 0.4185 0.4185 0.2657 0.4032 0.4009 0.3187 0.3152 0.0850 0.0144 0.4359
III-B1 Results on networks
@@
For the Dolphins, Polbooks, and Football datasets, each node is taken as the starting node to obtain the local community. For the Amazon and DBLP datasets, we randomly select 200 nodes from the top 5,000 communities as the starting nodes. The average values of Fscore and Jaccard for the starting nodes are shown in Table III.
Table III shows that ComGPT exhibits the best performance on the Amazon and DBLP datasets, ranking second and third places on the Dolphins and Polbooks datasets, respectively. On the Football dataset, ComGPT does not perform well. The reasons for the inconsistent performance of ComGPT across datasets are explained as follows. ComGPT achieves the highest performance on Amazon and DBLP. As shown in Table II, these datasets have low mixing parameters (0.1216 and 0.2147) and high modularity (0.44), indicating distinct community structures. This structure enables GPT-3.5-turbo to easily distinguish nodes inside and outside communities, thus achieving optimal results. For Polbooks and Dolphins, the modularity is lower than that of Amazon and DBLP, indicating weaker community structures. This increases the difficulty for GPT-3.5-turbo to accurately identify community members, resulting in some misidentified nodes and preventing ComGPT from achieving optimal performance. For Football, it has a high mixing parameter (0.3638), meaning numerous inter-community connections. These connections may cause GPT-3.5-turbo to merge multiple real communities into one, leading to ComGPT’s poor performance.
@@
LS performs well on datasets like Dolphins and Polbooks because it can accurately identify local leaders. This reliance on local leaders becomes a weakness in networks with strong inter-community connections, such as the football dataset. In these cases, a node identified as a local leader might actually be a follower of a more global leader, causing the algorithm to perform poorly. Louvain and its improved versions, such as Leiden and FLMIG, underperform on Amazon and DBLP datasets, as it tends to merge multiple distinct real communities into a single large community. Except for the DBLP dataset, DMF_M, DMF_R, and LCDNN perform better than the M method and R method. The reason is that the latter two methods maximize local modularity ($M$ and $R$), making them sensitive to the choice of seed nodes, known as seed-dependent problem. In contrast, DMF_M, DMF_R, and LCDNN implement various strategies to alleviate this problem. Specifically, DMF_M and DMF_R utilize membership functions to select nodes with large degrees, while LCDNN expands the community based on the fuzzy relations between nodes and their NGC nodes.
ComGPT aims to explore the application of LLMs to local community detection. The baselines are mature methods, optimized and tuned for excellent performance. In contrast, ComGPT, an emerging exploratory method based on LLMs, is still in its early stages and requires further optimization to reach its best performance. Its value lies in revealing the potential of LLMs in local community detection.
III-B2 Ablation study
(a) Dolphins
(b) Polbooks
Figure 6: Ablation study.
@@ III-B2 Ablation study
To explore the necessity of GPT-3.5-turbo, SK (Supplementary knowledge in graph coding), NSG, and NSU prompts in ComGPT, we conduct ablation experiments on the Dolphins and Polbooks datasets with 50 randomly selected seed nodes. Four simplified versions of ComGPT are constructed as follows: 1) Method 1 without GPT-3.5-turbo. The first node in $PN$ is selected for community expansion until $PN$ is empty; 2) Method 2 without SK; 3) Method 3 without NSG prompt; and 4) Method 4 without NSU prompt.
The simplified version of ComGPT may add duplicate nodes to the detected communities. When calculating Fscore and Jaccard, it is necessary to obtain the intersection between the detected community and the true community. Here, duplicate nodes in the detected community are regarded as a single node. Additionally, retain the duplicate nodes in the detected community. To prevent cyclic selection of duplicate nodes, community expansion is terminated once the number of duplicate nodes exceeds one-third of the number of nodes in the community.
@@
Figure 6 shows the results of Methods 1, 2, 3, 4, and ComGPT. Overall, ComGPT outperforms Methods 1, 2, 3, and 4, indicating that GPT-3.5-turbo, SK, NSG, and NSU prompts play a crucial role in ComGPT. ComGPT is better than Method 1 in terms of Fscore and Jaccard, which demonstrates the importance of GPT-3.5-turbo. The reason ComGPT outperforms Method 2 is that its graph coding incorporates community knowledge. Compared with graph encoding methods that contain only the graph topology, supplementing community knowledge enables the GPT-3.5-turbo to better understand the community structure and thus make more accurate judgments. ComGPT also outperforms Method 3 and 4, suggesting that NSG and NSU prompts can enhance its performance in local community detection.
III-B3 Prompt comparisons
TABLE IV: Description of the prompts.
@@ III-B3 Prompt comparisons
(a) Dolphins
(b) Polbooks
Figure 7: Results of different prompts.
To test the performance of the NSG prompt within ComGPT, we replaced the NSG prompt with four typical prompt approaches: Zero-shot [37], Few-shot [37], CoT [38], and BaG [22] prompts, as shown in Table IV. We conducted comparative experiments by randomly selecting 50 seed nodes in the Polbooks and Dolphins datasets. Detailed experimental results are shown in Figure 7. From these results, we derive the following conclusions:
Example samples are somewhat helpful to GPT-3.5-turbo. Based on the experimental results of Zero-shot and Few-shot, adding sample examples led to performance improvements in terms of Fscore and Jaccard on the Dolphins and Polbooks datasets. This indicates that providing examples for the GPT-3.5-turbo is indeed effective. Appropriate examples can improve performance.
Advanced prompts may have limited performance improvement or even counterproductive effects on complex graph reasoning tasks. For instance, advanced prompts like CoT can sometimes introduce more points of failure, as each step must be correct to ensure a successful final decision. It will increase the risk of error accumulation, especially in complex graph-based tasks. Consequently, advanced prompts may not always improve performance.
Compared to other prompts, NSG brings the highest improvement. Our designed prompt NSG achieve the best performance in terms of Fscore and Jaccard on Dolphins and Polbooks datasets. This is because the design of CoT and BaG is oriented towards universal graph reasoning problems so that no relevant guidance or knowledge is utilized in dealing with local community detection. NSG is oriented to the field of local community detection and utilizes the guidance in local community detection.
III-B4 Effect of NSG_1 and NSG_2 on ComGPT
We explore the effect of NSG_1 and NSG_2 on ComGPT on the polbooks dataset. In ComGPT, the NSG_1 prompt handles the seed-dependent problem and free rider effect, while the NSG_2 prompt tackles community diffusion (Section II-D). To verify whether ComGPT alleviates these shortcomings, we remove NSG_1 and NSG_2 from ComGPT to obtain methods named “No1” and “No2”, respectively. Result of No1, No2, ComGPT, and M Method on Polbooks is shown in Figure 8. Figure 8 shows that ComGPT, which combines the M method and GPT-3.5-turbo, outperforms the M method alone, indicating that ComGPT mitigates these shortcomings faced by the M method. ComGPT’s superiority over No1 indicates that the first prompt mitigates seed-dependent problem and free-rider effect, while its superiority over No2 shows that the second prompt mitigates community diffusion.
III-B5 Effect of the number of potential nodes on ComGPT
Figure 8: Effect of NSG_1 and NSG_2 on ComGPT.
@@ III-B5 Effect of the number of potential nodes on ComGPT
We explore the effect of the number of potential nodes, denoted by $k$, on ComGPT on the Football dataset, with 50 seed nodes selected randomly. The values of $k$ are 5, 7, 10, 13, and 15. Figure 9 shows the result of ComGPT under different $k$. As $k$ increases, the performance of ComGPT remains generally stable, though with some fluctuations. Fluctuations are relatively low across different values, except for $k$ = 10, which exhibits higher variability due to the randomness of GPT-3.5-turbo output. To confirm this, we repeat experiments for $k$ = 10 and the results are nearly identical to those obtained with other values of $k$. The above analysis suggests that the performance of ComGPT is less affected by the parameter $k$.
III-B6 Analysis of GPT-3.5-turbo API Runtime in ComGPT
TABLE V: Execution Times on node 1 of the Football dataset.
GPT-3.5-turbo Response Time Other Execution Time
t(s) Ratio(%) t(s) Ratio(%)
1 133.4220 99.9441 0.0745 0.0559
2 285.5551 99.9755 0.0699 0.0245
3 150.5789 99.9595 0.0609 0.0405
4 298.0469 99.9805 0.0579 0.0195
5 128.2163 99.9540 0.0589 0.0460
Average 199.1678 99.9677 0.0644 0.0323
To analyze the runtime of calling the GPT-3.5-turbo API in ComGPT, we selected node 1 of the Football network as the starting node. We record the execution time of ComGPT at this starting node over five runs, including the GPT-3.5-turbo response time, the ratio of GPT-3.5-turbo response time in ComGPT, the local computation time, the proportion of local computation time in ComGPT, and the average values of these results. The experimental results are shown in Table V.
The experimental results show that the response time of GPT-3.5-turbo accounts for more than 99.9% of the ComGPT’s execution time, indicating that the overall execution efficiency is mainly limited by GPT-3.5-turbo’s response speed rather than other computational processes. ComGPT needs to call GPT-3.5-turbo and wait for its response. This process is influenced by various factors, such as network conditions and server load, making it difficult to predict the exact response time.
III-B7 Scalability
To evaluate the scalability of ComGPT, we randomly sample subgraphs with different sizes from the Amazon dataset and test ComGPT on these extracted subgraphs. We extract subgraphs containing 20%, 40%, 60%, 80%, and 100% of the nodes from the Amazon dataset, respectively. We randomly select 100 nodes as the starting nodes from each subgraph. The average values of Precision, Recall, Fscore, Jaccard, runtime, and number of tokens are shown in Table VI.
Table VI shows that, as the percentage of nodes increases, the Recall, F-score, and Jaccard of ComGPT improve. This is because, as the number of nodes and edges in the network increases, the community structure becomes more obvious, and the communities detected by ComGPT gradually become more complete. Furthermore, as the percentage of nodes increases, we observe that the Precision fluctuates while the Recall consistently increases. This growth in Recall indicates that the size of the detected communities also increases with the network size, covering more ground-truth members. The fluctuation in Precision suggests that this expansion is not always perfectly accurate. The total runtime and number of tokens of ComGPT increase as the network grows. As the percentage of nodes increases, the average degree of the nodes in the network rises. Since the time complexity is proportional to the average degree, the runtime increases as the percentage of nodes increases. As the percentage of nodes increases, the size of the detected communities also increases, leading to a higher number of required API calls.
TABLE VI: Scalability of ComGPT on the Amazon dataset.
20% 40% 60% 80% 100%
Precision 0.9930 0.9657 0.9609 0.9520 0.9332
Recall 0.1497 0.3121 0.4779 0.6973 0.8601
Fscore 0.2452 0.4322 0.5954 0.7729 0.8646
Jaccard 0.1490 0.3038 0.4594 0.6719 0.8033
Time (s) 28.63 80.98 84.95 181.97 352.07
Tokens 383.14 1344.21 4026.44 11431.69 12386.43
IV Related work
In this section, we briefly review the related work on community detection and the application of Large Language Models (LLMs) in graph analysis.
IV-A Community detection
Community detection aims to identify groups of closely interconnected nodes. According to whether they utilize global information of the network, existing community detection methods are classified into two types: global community detection and local community detection.
IV-A1 Global community detection
Global community detection depends on the network’s global information to discover all communities. In recent years, researchers have investigated various algorithms [39], such as the Label Propagation Algorithm (LPA) [40], Non-negative Matrix Factorization (NMF) [41], random walk based techniques [42], and so on. Raghavan et al. introduced the Label Propagation Algorithm (LPA) [43]. In the LPA, each node iteratively updates its label to the label with the highest frequency. Communities are divided based on the distribution of labels. Wang et al. developed Non-negative Matrix Factorization (NMF)-based algorithms tailored for undirected networks, directed networks, and compound networks [44]. The algorithm determines the membership of nodes to communities by decomposing the adjacency matrix. Meo et al. proposed that edge weights enhance community detection and developed the WERW-Kpath [45]. It consists of two steps: (1) selecting a source node based on the degree of the node and the number of nodes in the network, and (2) choosing an unvisited edge from current node with probability proportional to its weight.
The above studies require complete network data to identify all communities. In contrast, our work focuses on determining the community of a specific node using only local information, thereby fundamentally distinguishing it from global studies.
IV-A2 Local community detection
Local community detection focuses on discovering the community to which the seed node belongs [46, 47], without involving the entire network. Compared to global community detection, it is more efficient and avoids unnecessary computational overhead. A variety of methods, like $k$-core decomposition [48] and seed expansion algorithms [24], are developed for detecting local communities. Seed expansion algorithms expand communities by selecting nodes with high scoring function values [9]. For example, the R method [24] and M method [4] choose nodes based on the local modularity R and M, respectively. The seed expansion algorithms often face the seed-dependent problem [10], community diffusion, and free rider effect [11, 12, 13]. To address the seed-dependent problem, researchers developed different strategies such as identifying core node sets [14, 49], finding the nearest nodes with greater centrality [15], alternating strategy of strong fusion and weak fusion [16], and designing membership functions with a bias towards nodes with large degrees [29]. Differing from the above studies, our work alleviates the aforementioned shortcomings by utilizing LLMs.
IV-B LLMs in graph analysis
IV-B1 Application of LLMs in graph reasoning
In recent years, researchers have begun to explore the potential of LLMs for applications in the areas of graph reasoning and graph machine learning. Wang et al. designed the NLGraph dataset for evaluating the graph reasoning capabilities of LLMs and demonstrated that LLMs do have graph reasoning abilities [22]. Zhang et al. designed the Graph-ToolFormer framework to handle graph reasoning tasks [50]. It enabled LLMs to self-guide using ChatGPT-enhanced prompts. Chen et al. utilized LLMs for the node classification task and proposed two approaches: LLMs-as-Enhancers and LLMs-as-Predictors [51]. The former leverages LLMs to enhance the node’s text attributes with their massive knowledge and then generates predictions through GNNs. The latter attempts to directly employ LLMs as standalone predictors. Chen et al. proposed a new label-free pipeline LLM-GNN to leverage LLMs for annotation, providing training signals on GNN for further prediction [52]. Specifically, LLMs are leveraged to annotate a small portion of nodes and then GNNs are trained on LLMs’ annotations to make predictions for the remaining large portion of nodes.
IV-B2 Graph encoding in LLMs
In the field of LLMs in graph machine learning research, scholars have focused on how to improve the ability of LLMs in processing graph data. Guo et al. emphasized the design and improvement of graph encoding techniques, which critically affect LLMs’ interpretation of input data and its subsequent output [53]. Therefore, Adjacency and Incident [23] are designed to encode graph structured data into text for LLMs. Adjacency uses integer node encoding and parenthesis edge encoding. Incident uses integer node encoding and incident edge encoding. Existing graph coding methods primarily represent graph topology, which is often insufficient for specific tasks, such as community detection. Unlike existing approaches, our work introduces a novel graph encoding method tailored for the community detection task.
IV-B3 Prompts in LLMs
PROMPT engineering is an approach to increase model inputs through prompts used for specific tasks so that LLMs are better adapted to new tasks [54]. The prompt is widely used in computer vision [55], recommender systems [51], and other fields. This approach does not require modification of model parameters, but rather significantly enhances the model’s ability to process downstream tasks based on task-specific prompts. Various prompts are designed for specific tasks to make LLMs better adapt to these tasks [54], such as the general CoT (Chain-of-Thought) prompt [38] and the BaG (Build a Graph) prompt [22]. CoT aims to enable LLMs to solve complex problems through step-by-step reasoning. Its core idea is to mimic human thought processes by reasoning incrementally rather than directly generating the final answer. BaG encourages LLMs to map the textual descriptions of graphs and structures to grounded conceptual spaces before tackling the specific problem through a one-sentence instruction. However, these prompts do not incorporate the domain knowledge required for local community detection and are unsuitable for detecting community structures within networks. This paper designs prompts suitable for local community detection tasks, which are different from the above prompts.
V Conclusion